OpenStudy (anonymous):

A(2,2) , B(-5,4) How would i find the slope of this, and whats the answer?

OpenStudy (anonymous):

Use A program Like Geogebra and it will make life easier, tust me'

OpenStudy (anonymous):

just use the point slope form

OpenStudy (anonymous):

do you know what point slope form is...the whole (y2-y1) / (x2-x1)

OpenStudy (anonymous):

roxgur1453 is also right ....i guess will ease calculations by a step or so

OpenStudy (anonymous):

yes, im just really confused on how to use the form.

OpenStudy (anonymous):

its really easy...so you have two ordered pairs. one orderd pair is x1 and y1, the other is x2 and y2. just put them into the equation then. so we'd have (4-2)/ (-5-2)

OpenStudy (anonymous):

so ordered pair A was x1, y1 and ordered pair B was x2, y2

OpenStudy (anonymous):

so would the answer be 2/-7?

OpenStudy (anonymous):

Yes!!! you got it!

OpenStudy (anonymous):

it helps to label your corrdinates, too. like write over to of them x1, y1 and x2, y2.

OpenStudy (anonymous):

okay...so what about, L(5,5) M(4,2)

OpenStudy (anonymous):

(5-2)/(5-4)

OpenStudy (anonymous):

try it and see what you get. label the points and just plug in the right x and y values

OpenStudy (anonymous):

okay. ill try. same way like we did last time?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

priceless6899 remember graphs will only come by practice .....try a lot of them and u can perfect them...

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

you get the answer yet?

OpenStudy (anonymous):

no not yet. kinda confused.

OpenStudy (anonymous):

no..i got 5-2, and 5-4.

OpenStudy (anonymous):

you're right... never mind haha. sorry i'm working on another probelm too!

OpenStudy (anonymous):

hahaha..yay! so did i get it right?

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

i did it like the opposite ya knoow....i used your 1st x&y as my second lol.no you should get 3

OpenStudy (anonymous):

i think you just subtracted wrong. you have the values right...5-2 = 3 and 5-4= 1. so it's be 3/1 not 1/3

OpenStudy (anonymous):

oh okay, i see what i did wrong! lol. how old r u?

OpenStudy (anonymous):

priceless try this out but now with a time limit of 2 mins :- (-30,-5) (-2,-3)

OpenStudy (anonymous):

A(2,2) , B(-5,4) Slope m = (4 -2)/ (-5 -2) = -2/7